### v2

@@ 3 The Italian participation in IUPAP in the interwar period
Volterra’s engagement was deeply intertwined with his and other Italian scientists’ ambitions to utilize the international institutional landscape as a means to reform the organization of scientific and technological research in Italy. In the lead-up to the establishment of the IRC, it was envisioned that each member country would set up a central state body to organize research on the model of the National Research Council, created in the United States in 1916.
@@
Volterra championed this project in Italy, initiating a process that would lead to the creation of the National Research Council (CNR, from its Italian name Consiglio Nazionale delle Ricerche), established in November 1923 within the Accademia dei Lincei, with Volterra being the chairman of both the CNR and Accademia dei Lincei [56]. Prior to the CNR’s inception, the Accademia dei Lincei served as Italy’s primary representative in international scientific fora. As President of the Accademia dei Lincei, Volterra played a pivotal role in orchestrating Italy’s involvement in IUPAP, contributing to the drafting of the preliminary statutes in 1919. He was also instrumental in nominating physicist Orso Mario Corbino, a fellow member of the Accademia dei Lincei, to the temporary ten-member IUPAP Executive Committee in July 1922.
@@
The interval between the establishment of IUPAP’s Executive Committee and its inaugural General Assembly in December 1923 coincided with dramatic transformations in Italian political landscape. The march on Rome by the National Fascist Party’s militants in October 1922 culminated in a regime change, with Benito Mussolini being appointed to form a new government. Although avowedly anti-fascist, Volterra initially collaborated with the new regime on matters concerning the country’s scientific and technological development and the creation of the CNR, which was finalized during the Mussolini government. The CNR’s bylaws mandated its affiliation with the IRC and called for the establishment of specialized commissions to integrate into the various unions under the IRC umbrella [20]. At the time of IUPAP’s formation, however, the CNR and these commissions had not been established yet. In fact, at the first IUPAP General Assembly Volterra was the only Italian representative. It wasn’t until December 1925, following the second IUPAP General Assembly and after the institutional limitations and the impact of national political changes on Italian engagement had become evident, that a physics committee within the CNR was constituted.
@@
The consolidation of the Fascist government into a dictatorship in 1925 precipitated the reformation of the CNR in 1927, leading to its detachment from the Accademia dei Lincei and its reconstitution as a governmental entity. In this new structure, the directorate was nominated by the government and reported directly to the Head of the Italian Government, namely, Benito Mussolini. This reform also resulted in the dismissal of Volterra from his roles within the CNR, with Guglielmo Marconi, who was more aligned with the Fascist regime, being installed as the CNR’s president. Despite this change, during the 1928 IRC assembly, Volterra’s position as vice-president of the IRC was reaffirmed. This decision faced opposition from the newly appointed Italian representatives of the CNR, who contested Volterra’s election on the grounds that he no longer served as an official Italian delegate, thus deeming his election illegal. [^11] Magrini to Arthur Schuster, September 5, 1928, Box 20, Folder 366, Fondo Consiglio Nazionale delle Ricerche, serie Presidenza Marconi (hereafter CNRMarconi), Archivio Centrale dello Stato, Rome (hereafter ACS); see also [45]. In response to this objection, the IRC Executive Committee staunchly defended the decision, asserting that Volterra’s election was fully compliant with IRC regulations. [^12] Arthur Schuster to Magrini, August 2, 1928, Box 20, Folder, 366, CNRMarconi.
@@
This scenario prompted Italy to adopt stringent positions against the IRC’s framework. In 1928, the IRC initiated discussions to revise its statutes, aiming for an organizational overhaul in 1931 that would potentially enable Germany and other previously excluded nations to join. The CNR advocated forcefully for the unions’ complete independence from the IRC, a stance largely driven by the desire to dissociate from anti-fascist sentiments perceived to be prevalent within the IRC’s leadership, as stressed by the vice-president of the CNR, jurist and diplomat Amedeo Giannini. [^13] A. Giannini to the secretary of the IRC, Sir H. Lyons, August 14, 1929, Box 20, Folder 367, CNRMarconi; “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche, 31 Ottobre 1930,” Archives of CNR, Roma (hereafter ACNR). By 1939, with the discussions on revising the IRC still underway, the CNR directorate started to evaluate Italian involvement in the individual unions. It solicited reports from the Italian commissions assessing the performance of various unions “because it is known that some unions work well, some mediocrely, and others do not work at all."141414“Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche,” October 31, 1929, ACNR, all translations from Italian documents are by the authors. Among the unions considered unnecessary by the directorate of the CNR there certainly was the physics Union, to which the CNR had not paid membership dues since 1927.
@@
The deliberations on Italian participation in the IRC became matter of foreign policy, of positioning Italy at the vanguard of an international campaign to boycott the IRC. It was concluded that “the [International Research Council] as it is organized serves no purpose," leading the directorate to recommend to Mussolini that Italy withdraw from the IRC for “essentially political" reasons. [^15] Promemoria, undated, Box 20, Folder 266, CNRMarconi; “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche” January 22, 1931; and “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche” February 13, 1931, ACNR. Beyond these political considerations, the directorate cited the proliferation of international organizations in which Italy held membership as overly burdensome. They suggested a more selective approach whereby Italy would engage with only one scientific organization per discipline, chosen with careful consideration of the organization’s quality and utility. It is not difficult to see that the CNR directorate did not regard IUPAP as meeting these criteria, viewing participation in such unuseful unions as not only unnecessary but also “harmful." [^16] Promemoria, undated, Box 20, Folder 266, CNRMarconi.
@@
This strategic decision was bolstered by a political campaign aimed at informing, through Italian ambassadors, all the countries that had previously been excluded or were still undecided about affiliating with the successor to the IRC, with a particular focus on Germany. [^17] Marconi to the Ministry of Foreign Affairs, February 5, 1931, Box 20, Folder 366, CNRMarconi; beyond Germany, the States informed by Italian position about the IRC membership were: Argentina, Austria, Bulgaria, Brazil, Hungary, Spain, Turkey. See, Sottosegretario di Stato to G. Marconi, October 14, 1931, Box 20, Folder 366, CNRMarconi. The goal was to publicly declare Italy’s stance on boycotting the IRC, thereby paving the way for other unaffiliated countries to sustain their dissociation from the IRC beyond 1931. This strategy aimed at potentially fostering a new framework for the organization of international science, with Italy assuming a leading role. Consequently, the transition from the IRC to ICSU in 1931 unfolded without Italy’s participation. And the same happened for IUPAP. At the 1931 General Assembly of the Union Italy was no longer among the member countries, while at the same time continuing its participation in the activities of other unions.181818“Union internationale de physique pure et appliquée. Troisième assemblée générale. Bruxelles, jullet 1931,” Serie B2aa, Vol. 1, IuG; for the continued participation in other unions, see sub-folder Mathematical Union, Box 20, Folder 367, Subfolder 8, CNRMarconi.
@@
These developments, combined with the impression that Germany was on the verge of joining ICSU and its unions, spurred a strategic reevaluation within the CNR directorate. In discussions with Mussolini, a consensus emerged that Italy should actively seek membership in ICSU and promptly inform all countries previously notified of Italy’s disengagement. [^20] G. Marconi to the Minister of Foreign Affairs, May 12, 1931, Box 20, Folder 366, CNRMarconi. After this decision, Italy rejoined IUPAP, even paying the back dues. This pivot resulted in Italy’s reintegration into IUPAP, inclusive of settling outstanding dues.212121“Union internationale de physique pure et appliquée. Quatrième assemblée générale. Londres, 5 october 1934,” Serie B2aa, Vol. 1, IuG. Consequently, Italy’s membership appeared uninterrupted, although our analysis reveals that, for a period, the CNR’s strategy was to maintain affiliations solely with selected unions deemed most beneficial, which did not include IUPAP. Italy’s engagement with IUPAP remained consistent up until the outbreak of World War II. However, IUPAP’s failure to formally incorporate German scientists and its subsequent inactivity led to a perception of Italy’s involvement as being in a state of perpetual stagnation, mirroring that of IUPAP itself. This view is further underscored by Enrico Fermi’s choice to decline the presidency of IUPAP following Niels Bohr’s refusal, signaling the organization’s unattractiveness to Italian physicists as a platform for international scientific collaboration.
@@
After World War II, IUPAP was reestablished on entirely new foundations. [^22] For a detailed discussion of IUPAP’s refoundation after World War II, see [33]. During the fifth IUPAP General Assembly in January 1947, nearly forty physicists convened in Paris to deliberate on the organization’s future direction. Reflecting on IUPAP’s shortcomings in the interwar period, the acting IUPAP Secretary General, UK-based German physicist Paul P. Ewald, underscored the necessity for a transformative approach in the organization’s activities. He advocated for principles of inclusivity and openness to all nations, including those that had been defeated during World War II “as soon as political conditions would permit."232323“IUPAP, Report 5th General Assembly, September 1947,” p. 17, Series B2aa, Vol. 1, IuG, translation by the authors.
@@
These institutional and individual refoundation efforts were significantly influenced by the geopolitical climate of the early Cold War era, a period marked by the political, ideological, economic, and military opposition between the Euro-Atlantic bloc and the Soviet bloc. In the early years of the Cold War, the Soviet Union pursued a policy of isolation, opting out of participation to international institutions, including those dedicated to scientific collaboration. Consequently, IUPAP predominantly consisted of members from the emerging Western bloc, supplemented by few non-aligned countries. As much as the principles accepted in 1947 declared an unconditional openness to all countries, the political landscape prevented an immediate embodiment of this ideal within IUPAP. Moreover, while the United States was officially a national member of IUPAP, U.S. physicists were notably absent from the first General Assembly. As the former President of IUPAP, Robert Millikan held an ex officio membership on the Union’s Executive Committee. He was therefore the liaison of the U.S. physics community with IUPAP. Ewald approached him to nominate U.S. representatives and delegates for the inaugural post-World War II General Assembly in Paris. [^24] Ewald to Millikan, August 7, 1946, Millikan Papers, Reel 12, p. 752. However, Millikan’s advanced age precluded his attendance, and the absence of a constituted U.S. national committee of physics meant that official U.S. delegates could not be determined in time for the meeting. [^25] Millikan to Fleury, February 28, 1947, Millikan Papers, Reel 12, p. 784; see also, IUPAP Circulaire d’Informtation générale, September 1947, Millikan Papers, Reel 12, p. 797. As a result, the fifth IUPAP General Assembly (the first to convene following the war) was de facto mostly a European affair. [^26] The only non-European physicists attending the meeting were representatives of China, and one representative from Australia. At the time both countries were Western allies, as the change of regime resulting from the victory of the Chinese Communist Party in mainland China will occur later, in 1949.
@@
The concurrent emergence of grassroots efforts by physicists to revitalize IUPAP and the development of a new institutional framework significantly bolstered the organization’s initiatives. Prior to this period, IUPAP had not established any commissions focused on particular research areas in physics, despite such provisions being included in its inaugural statutes. From the 1947 General Assembly onward, the formation of topical commissions emerged as a cornerstone of IUPAP’s activities. This shift signified a profound transformation for IUPAP: it transitioned from an entity primarily concerned with international standardization to an organization fostering international collaboration within distinct research fields of physics. Nowhere this change is most visible that in the creation of the initial topical commissions. Notably, the earliest of these was the commission that would evolve into the Commission on Thermodynamics and Statistical Mechanics. [^27] As discussed in the previous sections other ICSU unions were more active in the interwar period and had already established various topical commissions. Examples are the IAU and the International Union of Pure and Applied Chemistry (IUPAC), see [9, 18].

### v3

@@ 3 The Italian participation in IUPAP in the interwar period
Volterra, Senator of the Italian Kingdom since 1905, was one of the most active figures in organizing Italian science and promoting it abroad. His engagement with the IRC was deeply intertwined with his and other Italian scientists’ ambitions to utilize the international institutional landscape as a means to reform the organization of scientific and technological research in Italy. In the lead-up to the establishment of the IRC, it was envisioned that each member country would set up a central state body to organize research on the model of the National Research Council, created in the United States in 1916. Volterra championed this project in Italy, initiating a process that would lead to the creation of the National Research Council (CNR, from its Italian name Consiglio Nazionale delle Ricerche), established in November 1923 within the Accademia dei Lincei, with Volterra being the chairman of both the CNR and the Accademia dei Lincei [77]. Prior to the CNR’s inception, the Accademia dei Lincei served as Italy’s primary representative in international scientific fora. As president of the Accademia dei Lincei, Volterra played a pivotal role in orchestrating Italy’s involvement in IUPAP, contributing to the drafting of the preliminary statutes in 1919.131313Volterra’s personal commitment in the organization of the Italian physics community is attested by his status as one of the founding members of SIF in 1897, where he served as president from 1907 to 1909. He was also instrumental in nominating physicist Orso Mario Corbino, a fellow member of the Accademia dei Lincei and a Senator of the Kingdom, to the temporary ten-member IUPAP Executive Committee in July 1922.141414Corbino was also a leading figure in the revival of Italian physics in the late 1920s. One of his major accomplishments was recognizing that Enrico Fermi could play a key role in spreading modern physics in Italy and educating new generations. With support from influential mathematicians like Volterra, Corbino facilitated the establishment of Italy’s first chair of theoretical physics, to which Fermi was appointed in 1926, see [28].
@@
The interval between the establishment of IUPAP’s Executive Committee and its inaugural General Assembly in December 1923 coincided with dramatic transformations in the Italian political landscape. The march on Rome by the National Fascist Party’s militants in October 1922 culminated in a regime change, with Benito Mussolini being appointed to form a new government. Although avowedly anti-fascist, Volterra initially collaborated with the new regime on matters concerning the country’s scientific and technological development and the creation of the CNR, which was finalized during the Mussolini government. The CNR’s bylaws mandated its affiliation with the IRC and called for the establishment of specialized commissions to integrate into the various unions under the IRC umbrella [32]. At the time of IUPAP’s formation, however, the CNR and these commissions had not been established yet. In fact, at the first IUPAP General Assembly Volterra was the only Italian representative. It wasn’t until December 1925, following the second IUPAP General Assembly and after the institutional limitations and the impact of national political changes on Italian engagement had become evident, that a physics committee within the CNR was constituted.
@@
The consolidation of the Fascist government into a dictatorship in 1925 precipitated the reformation of the CNR in 1927, leading to its detachment from the Accademia dei Lincei and its reconstitution as a governmental entity. In this new structure, the directorate was nominated by the government and reported directly to the Head of the Italian Government, namely, Benito Mussolini. This reform also resulted in the dismissal of Volterra from his roles within the CNR, with Guglielmo Marconi, who was more aligned with the Fascist regime, being installed as the CNR’s president. [^15] Later, after Volterra refused to take the oath of loyalty to Mussolini’s government in 1931, he was forced to leave the University of Rome and resign from all Italian scientific academies, where he had long been one of the most distinguished members. Despite this change, during the 1928 IRC assembly, Volterra’s position as vice-president of the IRC was reaffirmed. This decision faced opposition from the newly appointed Italian representatives of the CNR, who contested Volterra’s election on the grounds that he no longer served as an official Italian delegate, thus deeming his election illegal. [^16] Magrini to Arthur Schuster, September 5, 1928, Box 20, Folder 366, Fondo Consiglio Nazionale delle Ricerche, serie Presidenza Marconi (hereafter CNRMarconi), Archivio Centrale dello Stato, Rome (hereafter ACS); see also [63]. In response to this objection, the IRC Executive Committee staunchly defended the decision, asserting that Volterra’s election was fully compliant with IRC regulations. [^17] Arthur Schuster to Magrini, August 2, 1928, Box 20, Folder, 366, CNRMarconi.
@@
This scenario prompted Italy to adopt stringent positions against the IRC’s framework. In 1928, the IRC initiated discussions to revise its statutes, aiming for an organizational overhaul in 1931 that would potentially enable Germany and other previously excluded nations to join. The CNR advocated forcefully for the unions’ complete independence from the IRC, a stance largely driven by the desire to dissociate from anti-fascist sentiments perceived to be prevalent within the IRC’s leadership, as stressed by the vice-president of the CNR, jurist and diplomat Amedeo Giannini. [^18] A. Giannini to the secretary of the IRC, Sir H. Lyons, August 14, 1929, Box 20, Folder 367, CNRMarconi; “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche, 31 Ottobre 1930,” Archives of CNR, Roma (hereafter ACNR). By 1939, with the discussions on revising the IRC still underway, the CNR directorate started to evaluate the Italian involvement in the individual unions. It solicited reports from the Italian commissions assessing the performance of various unions “because it is known that some unions work well, some mediocrely, and others do not work at all."191919“Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche,” October 31, 1929, ACNR, all translations from Italian documents are by the authors. Among the unions considered unnecessary by the directorate of the CNR there certainly was IUPAP, to which the CNR had not paid membership dues since 1927.
@@
The deliberations on Italian participation in the IRC became matter of Italian foreign policy, with the aim of positioning Italy at the vanguard of an international campaign to boycott the IRC. It was concluded that “the [International Research Council] as it is organized serves no purpose," leading the directorate to recommend to Mussolini that Italy withdraw from the IRC for “essentially political" reasons. [^20] Promemoria, undated, Box 20, Folder 266, CNRMarconi; “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche,” January 22, 1931; and “Verbale della seduta del direttorio del Consiglio Nazionale delle Ricerche,” February 13, 1931, ACNR. Beyond these political considerations, the directorate cited the proliferation of international organizations in which Italy held membership as overly burdensome. They suggested a more selective approach whereby Italy would engage with only one scientific organization per discipline, chosen with careful consideration of the organization’s quality and utility. It is not difficult to see that the CNR directorate did not regard IUPAP as meeting these criteria, viewing participation in such unuseful unions as not only unnecessary but also “harmful." [^21] Promemoria, undated, Box 20, Folder 266, CNRMarconi.
@@
This strategic decision was bolstered by a political campaign aimed at informing, through Italian ambassadors, all the countries that had previously been excluded or were still undecided about affiliating with the successor to the IRC, with a particular focus on Germany. [^22] Marconi to the Ministry of Foreign Affairs, February 5, 1931, Box 20, Folder 366, CNRMarconi; beyond Germany, the States informed of the Italian position about the IRC membership were: Argentina, Austria, Bulgaria, Brazil, Hungary, Spain, Turkey. See, Sottosegretario di Stato to G. Marconi, October 14, 1931, Box 20, Folder 366, CNRMarconi. The goal was to publicly declare Italy’s stance on boycotting the IRC, thereby paving the way for other unaffiliated countries to sustain their dissociation from the IRC beyond 1931. This strategy aimed at potentially fostering a new framework for the organization of international science, with Italy assuming a leading role. Consequently, the transition from the IRC to ICSU in 1931 unfolded without Italy’s participation. And the same happened for IUPAP. At the 1931 General Assembly of the Union Italy was no longer among the member countries, while at the same time continuing its participation in the activities of other unions.232323“Union internationale de physique pure et appliquée. Troisième assemblée générale. Bruxelles, juillet 1931,” Serie B2aa, Vol. 1, IuG; for the continued participation in other unions, see sub-folder Mathematical Union, Box 20, Folder 367, Sub-folder 8, CNRMarconi.
@@
These developments, combined with the impression that Germany was on the verge of joining ICSU and its unions, spurred a strategic reevaluation within the CNR directorate. In discussions with Mussolini, a consensus emerged that Italy should actively seek membership in ICSU and promptly inform all countries previously notified of Italy’s disengagement. [^25] G. Marconi to the Minister of Foreign Affairs, May 12, 1931, Box 20, Folder 366, CNRMarconi. After this decision, Italy rejoined IUPAP, even paying the back dues. This pivot resulted in Italy’s reintegration into IUPAP, inclusive of settling outstanding dues.262626“Union internationale de physique pure et appliquée. Quatrième assemblée générale. Londres, 5 octobre 1934,” Serie B2aa, Vol. 1, IuG. Consequently, Italy’s membership appeared uninterrupted, although our analysis reveals that, for a period, the CNR’s strategy was to maintain affiliations solely with selected unions deemed most beneficial, which did not include IUPAP. Italy’s engagement with IUPAP remained consistent up until the outbreak of World War II. However, IUPAP’s failure to formally incorporate German scientists and its subsequent inactivity led to a perception of Italy’s involvement as being in a state of perpetual stagnation, mirroring that of IUPAP itself. This view is further highlighted by Enrico Fermi’s choice to decline the presidency of IUPAP following Niels Bohr’s refusal, signaling the organization’s unattractiveness to Italian physicists as a platform for international scientific collaboration.
@@
After World War II, IUPAP was reestablished on entirely new foundations. [^27] For a detailed discussion of IUPAP’s refoundation after World War II, see [49]. During the fifth IUPAP General Assembly in January 1947, nearly forty physicists convened in Paris to deliberate on the organization’s future direction. Reflecting on IUPAP’s shortcomings in the interwar period, the acting IUPAP secretary general, UK-based German physicist Paul P. Ewald, underlined the necessity for a transformative approach in the organization’s activities. He advocated for principles of inclusivity and openness to all nations, including those that had been defeated during World War II “as soon as political conditions would permit."282828“IUPAP, Report 5th General Assembly, September 1947,” p. 17, Series B2aa, Vol. 1, IuG, translation by the authors.
@@
These institutional and individual refoundation efforts were significantly influenced by the geopolitical climate of the early Cold War era, a period marked by the political, ideological, economic, and military opposition between the Euro-Atlantic bloc and the Soviet bloc. In the early years of the Cold War, the Soviet Union pursued a policy of isolation, opting out of participation in international institutions, including those dedicated to scientific collaboration. Consequently, IUPAP predominantly consisted of members from the emerging Western bloc, supplemented by few non-aligned countries. As much as the principles accepted in 1947 declared an unconditional openness to all countries, the political landscape prevented an immediate embodiment of this ideal within IUPAP. Moreover, while the United States was officially a national member of IUPAP, U.S. physicists were remarkably absent from the first General Assembly. As the former president of IUPAP, Robert Millikan held an ex officio membership on the Union’s Executive Committee. He was therefore the liaison of the U.S. physics community with IUPAP. Ewald approached him to nominate U.S. representatives and delegates for the inaugural post-World War II General Assembly in Paris. [^29] Ewald to Millikan, August 7, 1946, Millikan Papers, Reel 12, p. 752. However, Millikan’s advanced age precluded his attendance, and the absence of a constituted U.S. national committee of physics meant that official U.S. delegates could not be determined in time for the meeting. [^30] Millikan to Fleury, February 28, 1947, Millikan Papers, Reel 12, p. 784; see also, IUPAP Circulaire d’Information générale, September 1947, Millikan Papers, Reel 12, p. 797. As a result, the fifth IUPAP General Assembly (the first to convene following the war) was de facto mostly a European affair. [^31] The only non-European physicists attending the meeting were representatives of China, and one representative from Australia. At the time both countries were Western allies, as the change of regime resulting from the victory of the Chinese Communist Party in mainland China would occur later, in 1949.
@@
The concurrent emergence of grassroots efforts by physicists to revitalize IUPAP and the development of a new institutional framework significantly bolstered the organization’s initiatives. Prior to this period, IUPAP had not established any commissions focused on particular research areas in physics, although such provisions were included in its founding statutes. From the 1947 General Assembly onward, the formation of topical commissions emerged as a cornerstone of IUPAP’s activities. This shift signified a profound transformation for IUPAP: it transitioned from an entity primarily concerned with international standardization to an organization fostering international collaboration within distinct research fields of physics. Nowhere this change is most visible that in the creation of the initial topical commissions. Notably, the earliest of these was the commission that would evolve into the Commission on Thermodynamics and Statistical Mechanics. [^32] As discussed in the previous sections, other ICSU unions were more active in the interwar period and had already established various topical commissions. Examples are the IAU and the International Union of Pure and Applied Chemistry (IUPAC), see [7, 26].
In the wake of the 1947 General Assembly, the proliferation of commissions marked a significant phase of growth for IUPAP, but the scope and function of these commissions initially lacked clarity and differed from their eventual roles. Prior to 1947, IUPAP’s structure mainly consisted of general-purpose commissions, such as the SUN Commission and the Commission on Publications. The 1947 General Assembly saw the establishment of several new commissions alongside the reconstitution of the SUN Commission, which was inactive since the mid-1930s (see Table 1).